# Graph-Based Fault Diagnosis for Rotating Machinery: Adaptive Segmentation and Structural Feature Integration

### v1

@@ 4.3.4 Feature Importance Analysis
To elucidate the predictive behavior of the logistic regression model, we analyze its learned coefficients, where positive coefficients enhance the likelihood of the positive class (fault conditions), and negative coefficients align with the negative class (healthy conditions). This analysis, illustrated in Figure 6, reveals features’ relative importance and direction at 0HP, 1Hp, 2 HP and 3 HP load. PSD Mid Freq Ch1 notably (for example in 2 Hp load) exhibits the highest positive coefficient ($\approx+1.9$), underscoring its critical role in fault prediction. Other influential positive features include PSD High Freq Ch1 and FFT High Freq Ch2, while features like FFT Mean Ch2, Skew Ch1, and RMS Ch2 display strong negative coefficients, likely reflecting baseline or healthy states.
Figure 6: Logistic regression coefficients showing the relative importance and direction of each feature for 0HP, 1 HP, 2 HP and 3 Hp load.
Here, Ch1 and Ch2 denote Channel 1 and 2, respectively, representing distinct vibration signal acquisition channels from the monitored system. The analysis indicates that Channel 1 features are more predictive of faults, whereas Channel 2 features correlate with regular operation. Frequency-domain features, such as PSD and FFT, dominate across both channels, with time-domain statistical measures (e.g., skewness, RMS) also contributing, albeit with channel-specific variations.
Graph-theoretic properties, including Modularity (Q), Spectral Gap, and Average Path Length, exhibit lower but non-negligible coefficients. While not primary drivers in the linear model, these features capture structural signal characteristics that aid classification.
To further investigate the contribution of features in the SU gearbox and bearing datasets, Figure 7 presents a heatmap of learned coefficients for the gearset/bearing data. This visualization highlights the influence of time- and frequency-domain features across various sensor channels. Warm colors (red) indicate a positive contribution to fault classification, while cool colors (blue) represent negative contributions. In the gearbox dataset, frequency-domain features such as FFT High Freq and FFT Low Freq in Channels 4 and 5 show strong positive influence, underscoring their importance. Time-domain statistical features like Mean, Median, and Kurtosis from the earlier channels also demonstrate notable relevance. Conversely, Channels 6 to 8 exhibit mostly neutral weights, indicating a relatively lower impact. A similar heatmap can be generated for the bearing dataset to analyze the influence of different features on classification decisions
Figure 7: Heatmap of logistic regression coefficients for the SU gearset and bearing dataset, illustrating the contribution of time- and frequency-domain features across sensor channels. Red shades indicate a positive influence on classification, while blue shades denote negative contributions.
This analysis highlights the interpretability of logistic regression, guiding effective feature selection for fault classification. Extending this approach to other datasets can ensure consistency and generalizability across diverse operating conditions.
4.3.5 Sensitivity Analysis
Figure 8: Classification accuracy under increasing levels of feature noise across different operating conditions. The top row shows performance on the CWRU bearing dataset under four machine load settings (0 HP, 1 HP, 2 HP, and 3 HP). The bottom row shows results on the SU Bearing and SU Gearbox datasets. As noise increases, accuracy declines, with higher loads exhibiting greater sensitivity to noise. In contrast, the SU datasets demonstrate stronger robustness across all noise levels.
This study further investigates the robustness of machine learning models to feature noise using three distinct vibration datasets: the CWRU bearing dataset, evaluated under varying machine loads (0 HP to 3 HP), and the SU Bearing and SU Gearbox datasets, assessed under standard operating conditions, as illustrated in Figure 8. In each case, Gaussian noise with increasing standard deviations (0, 0.05, 0.1, 0.2, and 0.5) is added to the input features, and its impact on classification accuracy is evaluated through k-fold cross-validation.
For the CWRU dataset, model performance remains stable under light noise at all load levels, but accuracy steadily declines as noise increases. At 0 HP, the model maintains high accuracy (approximately 0.982) up to a noise level of 0.2, after which it drops moderately to around 0.955 at 0.5 noise. As the load increases to 1 HP and 2 HP, the model becomes more sensitive, with accuracy decreasing to around 0.89 and 0.86, respectively, at the highest noise level. The most significant degradation is observed at 3 HP, where accuracy falls from approximately 0.982 to below 0.88 under maximum noise, indicating that higher operational loads exacerbate the model’s sensitivity to noisy input features.
In contrast, the SU datasets exhibit strong robustness to noise. The model starts with very high baseline accuracy (approximately 0.992 and 0.994, respectively) and maintains excellent performance up to moderate noise levels (0.2 standard deviation) for both the SU Bearing and SU Gearbox datasets. Even at the highest noise level (0.5), accuracy only dips slightly—to around 0.982 for SU Bearing and 0.979 for SU Gearbox—highlighting their stability and noise resilience. These findings suggest that while noise has a more pronounced effect under varying load conditions in the CWRU dataset, the models trained on the SU datasets remain consistently robust, making them highly reliable for real-world fault diagnosis, even in noisy signals.
@@ 4.4 Comparative Analysis
The proposed bearing fault detection model outperforms deep learning approaches (e.g., CNN, LSTM) and graph-based methods (e.g., GCN, GAT) on the CWRU and SU datasets. By integrating feature engineering, signal processing, and graph metrics (e.g., path length $L$, modularity $Q$, spectral gap) with logistic regression, the model achieves superior accuracy, computational efficiency, and interpretability. Performance benchmarks in Table 8 (accuracy) and Table 9 (inference time) demonstrate its clear advantage over more complex alternatives.
Table 8: Comparative Performance of Bearing Fault Detection Models on CWRU and PU Datasets
@@
WDCNN (Gao et al., 2024) 84.57
MFFCNN (Gao et al., 2024) 86.47
GAE (Singh, 2025) 97.6
Proposed Model 99.89 99.99
@@
Table 9: Inference Time Comparison on the CWRU and SU Dataset
@@
The model transforms vibrational data into a $(20ch+3)$-dimensional feature vector, where $C$ denotes the number of channels. For each channel, it extracts 8 time-domain features (e.g., mean, RMS, skewness), 12 frequency-domain features (e.g., dominant FFT frequency, PSD peak), and 3 graph-based metrics, resulting in 43 features when $ch=2$ (as in the CWRU and GearBox datasets). Features such as FFT Dom Freq Ch1 capture fault harmonics (e.g., 50–100 Hz on CWRU, 100–150 Hz on SU), while graph metrics like $Q$ quantify structural clustering. When trained and tested on the same dataset without cross-dataset learning, logistic regression achieves 99.89% accuracy on CWRU and 100.00% on SU (Table 8), outperforming more complex deep learning models and graph convolutional networks, which often suffer from heavy preprocessing requirements and limited interpretability.
On CWRU, the model’s 99.89% accuracy is compatible with MBH-LPQ-VGGish (98.95%) (Chennana et al., 2025), out performs DenseNet-12 (98.57%) (Niyongabo et al., 2022), and WDCNN (84.57%) (Gao et al., 2024). On SU, it achieves 100.00%, surpassing SCNN 3D (99.93%) (Xu et al., 2025) and Sparse Transformer (99.98%) (Zhou et al., 2024).
The model’s inference time is 0.000252 s on CWRU and 0.0001639 s on SU (Table 9), significantly faster than CBAM-MFFCNN (0.1179 s) (Gao et al., 2024), VGGish (1.769 s) (Chennana et al., 2025), MBH-LPQ-VGGish (1.998 s)  (Chennana et al., 2025), SCNN (0.002 s) (Xu et al., 2025), and DL-CNN (536 s) (Shao et al., 2018). Precomputed features reduce complexity, enabling real-time fault detection, unlike resource-intensive deep learning and GCNs.
SHAP analysis enhances the model’s interpretability. For eaxmple, the feature FFT Dom Freq Ch1 (SHAP: -15 to +15 on SU gearbox dataset) links to fault frequencies, and $Q$ (SHAP: 2–4) captures clustering. Similar trends appear on CWRU (SHAP: -20 to +20). Unlike deep learning and GCNs, which rely on vague post-hoc methods like Grad-CAM, the model’s features directly connect to physical phenomena, simplifying diagnosis.
The model performs well with limited data, achieving 99.89% on CWRU , 99.99% on SU with small subgraphs ($|\mathcal{G}_{i}|\leq 200$). Engineered features ensure strong generalization, unlike LSTMs (F1: 0.90 on small data) or GCNs, which falter without large graph samples.
In transfer learning experiments, the direction of training plays a critical role in determining generalization performance. On the CWRU dataset, training the model on 1 HP load data and testing across other load conditions (0, 2, and 3 HP) yields a higher average accuracy (99.00%) than training on 3 HP data (98.63%). This indicates that features learned under moderate loading conditions (1 HP) are more transferable, likely because they encapsulate characteristics of both low and high load dynamics. A similar trend is observed on the SU dataset, where training on Gearbox Gear data results in better transfer accuracy (97.3%) when tested on Bearing Gear data, compared to the reverse direction (94.9%). Notably, existing methods do not perform true transfer learning; instead, they train and test on mixed data from each vibrational dataset. This practice can lead to inflated performance metrics and limits the evaluation of a model’s ability to generalize across different domains or operating conditions (Zhao et al., 2019).
4.5 Discussion
The results from both the CWRU and SU datasets demonstrate the effectiveness and adaptability of the proposed graph-based fault detection approach across a wide range of operational scenarios and fault conditions. One of the most significant observations is the model’s strong performance across varying machine loads, particularly in the CWRU dataset. While accuracy and F1-scores were consistently high across all load conditions, classification performance slightly dipped under low-load scenarios (e.g., 0 HP), where vibration signals are less distinguishable due to weaker fault-induced signatures. In contrast, higher loads (2 HP and 3 HP) produced more pronounced fault features, contributing to improved classification metrics and better graph separability.
In the SU datasets, where operational settings were more controlled and balanced across fault types, the model demonstrated nearly perfect classification across all fault classes. These results highlight the robustness of the proposed method in identifying subtle vibration pattern changes, even across datasets with differing domain characteristics.
Additionally, the transfer learning experiments revealed that while same-load source–target configurations led to near-perfect classification, cross-load performance remained strong but showed slight drops depending on the disparity in load conditions. Statistical analysis validated that these variations were significant, reinforcing the need for adaptive models or fine-tuning in real-world, variable-load deployments.
Another key finding relates to graph construction parameters. Adjusting $k_{max}$ and subgraph sizes significantly influenced model efficiency and structural clarity. Small-world characteristics and higher modularity scores were more pronounced in mid-sized subgraphs, balancing detection performance with computational efficiency. Feature noise analysis further confirmed the model’s resilience, especially in the SU datasets, where accuracy remained stable even under high noise levels.

### v2

@@ 3.3.5 Feature Importance Analysis
To evaluate the Random Forest (RF) model’s predictive behavior under varying loads, SHAP (SHapley Additive exPlanations) analysis quantifies feature contributions in the CWRU and SU datasets. The CWRU dataset includes vibration signals from two channels (Ch1, Ch2), while the SU dataset comprises eight channels (Ch1–Ch8).
Table 11: Summary of SHAP-Based Feature Importance Across Load Conditions for CWRU dataset
Rank Feature Domain Ch. 0 Hp 3 Hp Trend
1 Mean Ch1 Time Ch1 Very High Low ↓ Decreasing
2 PSD Low Freq Ch1 Frequency Ch1 Very High Very High ↑ Then Stable
3 FFT Low Freq Ch2 Frequency Ch2 High Medium Stable / ↓
4 Kurtosis Ch1/Ch2 Time Ch1/2 High Low ↓ Decreasing
5 PSD Mid Freq Ch2 Frequency Ch2 Medium Very High ↑ Increasing
6 PSD High Freq Ch2 Frequency Ch2 Low Very High ↑ Increasing
7 TKE Ch1 Time-Frequency Ch1 Low Medium ↑ Then Stable
8 RMS Ch1 Time Ch1 Medium Low ↓ Slightly
9 FFT Std Ch1 Frequency Ch1 Low Medium Stable
10 Env FFT Mean Ch2 Envelope Freq Ch2 Medium High ↑ Increasing
11 FFT Mean Ch2 Frequency Ch2 Low Medium Stable
12 PSD Peak Ch1/Ch2 Frequency Ch1/2 Low Medium Stable
13 Median Ch2 Time Ch2 High Low ↓ Disappears
(a) SHAP feature importance at 0 HP (no load).
(b) SHAP feature importance at 3 HP (full load).
(c) SHAP Feature Importance – Bearing.
(d) SHAP Feature Importance – Gearbox.
Figure 9: Comparison of SHAP feature importance across CWRU and SU Datasets.
Table 12: Comparison of SHAP-Based Feature Importance Between SU Bearing and Gearbox Fault Classification
Rank Feature Domain Bearing Impact Gearbox Impact Trend
1 Median Ch1 Time Very High Medium ↓ Less Important in Gearbox
2 PSD Peak Ch5 Frequency High Very High ↑ More Dominant in Gearbox
3 Std Ch5 Time Very High Very High → Stable Across Both
4 Mean Ch1 Time Very High High ↓ Slight Drop in Gearbox
5 FFT Std Ch1 Frequency High Medium ↓ Lower in Gearbox
6 RMS Ch1 Time Medium High ↑ Higher in Gearbox
7 PSD Sum Ch1 Frequency Medium Medium → Similar Impact
8 RMS Ch5 Time Medium Medium → Stable
9 PSD Low Freq Ch5 Frequency High Medium ↓ Slight Drop
10 FFT Low Freq Ch1 Frequency High Medium ↓ Slight Drop
11 FFT Std Ch5 Frequency Medium Medium → Similar Impact
12 TKE Ch2 Time-Frequency Medium Low ↓ Less in Gearbox
13 Env FFT Mean Ch2 Envelope Freq Medium Low ↓ Declining
14 PSD Low Freq Ch3 Frequency Low Medium ↑ Only Important in Gearbox
15 FFT High Freq Ch4 Frequency Medium Low ↓ Less Important in Gearbox
The SHAP analysis results are summarized in Table 11 and detailed through visualizations in Figures 9(a) and 9(b), which illustrate the magnitude and direction of feature influences under the two load extremes. Under no-load conditions (0 HP), time-domain features like Mean Ch1 and Kurtosis Ch1/Ch2 strongly contribute to model predictions, reflecting their sensitivity to low-stress scenarios. However, their influence diminishes significantly under full-load conditions (3 HP). In contrast, frequency-domain features, such as PSD Mid Frequency Ch2, PSD High Frequency Ch2, and Envelope FFT Mean Ch2, become more prominent at full load, underscoring their importance in high-stress environments. Features like FFT Standard Deviation Ch1 and PSD Peak Ch1/Ch2 maintain relatively stable contributions across both load levels, indicating robustness to changing operational conditions.
Table12 and Figures9(c) and 9(d) compare SHAP feature importance for bearing and gearbox fault classification in the SU dataset. For bearings, time-domain features (Median Ch1, Std Ch5, Mean Ch1) dominate, capturing amplitude variations from localized faults (e.g., inner race defects). For gearboxes, frequency-domain features (PSD Peak Ch5, FFT Std Ch1) are critical, detecting spectral patterns linked to gear faults (e.g., root cracks). Features like Std Ch5 and RMS Ch1 maintain high importance across both, while TKE Ch2 and Env FFT Mean Ch2 are bearing-specific. Thus, time-domain features drive bearing diagnostics, while frequency-domain features are key for gearboxes.
3.3.6 Noise Sensitivity Analysis
(a) Accuracy with feature noise (0 Hp data load)
(b) Accuracy with feature noise (1 Hp data load)
(c) Accuracy with feature noise (2 Hp data load)
(d) Accuracy with feature noise (1 Hp data load)
(e) Accuracy with feature noise (SU Gearbox)
(f) Accuracy with feature noise (SU Bearing)
Figure 10: Classification accuracy of Random Forest under varying levels of Gaussian noise across datasets.
Random Forest demonstrates robust performance against noisy inputs due to its ensemble structure and randomized feature selection. This study evaluates RF’s resilience under Gaussian noise with zero-mean and standard deviations of $\sigma=0,0.05,0.1,0.2,0.5$ applied to input features of the CWRU bearing dataset (0–3 HP loads) and SU gearbox/bearing dataset. Classification accuracy, assessed via stratified 5-fold cross-validation, is reported in Figure 10.
For the SU dataset (Figure 10(e), and 10(f)), RF maintains high accuracy (0.994–0.980 and 0.992–0.982, respectively) across all noise levels, with a minimal decline of 1.2–1.4% at $\sigma=0.5$.
For the CWRU dataset at 1–3 HP (Figures 10(b) to 10(d)), accuracy remains robust at 0.99–0.98 for $\sigma\leq 0.1$, but drops to 0.88–0.84 at $\sigma=0.5$, indicating higher noise sensitivity in higher-load conditions. In contrast, at 0 HP (Figure 10(a)), accuracy is more stable, declining from 0.98 at $\sigma=0.05$ to 0.96 at $\sigma=0.5$, suggesting lower noise sensitivity in low-load conditions. This robustness, consistent with Breiman (Breiman, 2001), stems from RF’s ensemble averaging, where decision trees trained on bootstrapped samples and random feature subsets mitigate the impact of noisy features. The increased noise sensitivity at 1–3 HP may reflect greater signal complexity in higher-load vibration data.
These findings affirm RF’s suitability for fault diagnosis in low-to-moderate noise industrial settings, particularly for gearbox and low-load bearing applications. However, preprocessing or feature engineering may be required for higher-load scenarios under high noise ($\sigma\geq 0.5$).
@@ 3.4 Comparative Analysis
The proposed bearing fault detection model consistently outperforms conventional deep learning approaches (e.g., CNN, LSTM) as well as advanced graph-based techniques (e.g., GCN, GAT) on both the CWRU and SU datasets. By integrating multiscale feature engineering, signal processing, and graph-theoretic metrics—such as path length ($L$), modularity ($Q$), and spectral gap—within the random forest framework, the model achieves a compelling balance of accuracy, computational efficiency, and interpretability. Performance benchmarks in Table 13 (classification accuracy) and Table 14 (inference time) clearly illustrate the superiority of the proposed method over more complex and computationally intensive alternatives.
Specifically, Table 13 presents a comparative analysis of the proposed model against a wide range of state-of-the-art fault diagnosis techniques. The proposed approach attains exceptional classification accuracies of 99.89% on the CWRU dataset and 99.99% on the SU dataset, surpassing traditional CNN-based architectures (e.g., VGG-16, DenseNet-12, Inception-LSTM), graph-based models (e.g., GATLSTM, GNNBFD, BFDGE), and recent transformer-based solutions (e.g., Sparse Transformer). While some models, such as SCNN-3D (99.93%) and Sparse Transformer (99.98%), perform competitively on SU, they do not maintain equally high accuracy across both datasets. In contrast, the proposed model delivers consistently top-tier performance and demonstrates strong generalizability under varying load and fault conditions, establishing a new benchmark in fault diagnosis accuracy and efficiency.
Table 13: Comparative performance of bearing fault detection models on the CWRU and SU datasets. Accuracy values for baseline models are adopted from the cited sources; proposed model results are from our experiments.
@@
WDCNN (Gao et al., 2024a) 84.57
MFFCNN (Gao et al., 2024a) 86.47
GAE (Singh, 2025) 97.60
GATLSTM (Singh et al., 2024) 96.50
GNNBFD (Xiao et al., 2023) 98.03
BFDGE (Wang et al., 2024) 97.66
Proposed Model 99.90 100.00
@@
Note: Accuracy values for baseline models are reported directly from the respective cited works, based on experiments using the same datasets and conditions. No re-implementation of these models was performed.
Table 14: Inference Time Comparison on the CWRU and SU Dataset. Source: Authors own work.
@@
Table 15: Computational performance metrics across experimental conditions
Condition Segmentation Feature Extraction Per-Window Graph Construction
(s/instance) (s/instance) (s/window) (s/instance)
1 HP Load (CWRU) 0.0003 1.0613 0.0050 0.0997
2 HP Load (CWRU) 0.0005 1.1283 0.0032 0.1377
3 HP Load (CWRU) 0.0004 1.3435 0.0039 0.1038
0 HP Load (CWRU) 0.0001 0.1940 0.0033 0.0293
Gearset (SU) 0.0013 17.7302 0.0099 0.3376
BearingBox (SU) 0.0015 23.0519 0.0108 0.3285
[…]
To provide a finer breakdown of computational cost, Table 15 details the per-instance execution time across key processing stages of segmentation, feature extraction, window-level inference, and graph construction under various load conditions. The proposed model maintains highly efficient performance despite incorporating graph-based reasoning, which is typically computationally intensive. For example, total feature extraction time remains under 1.35 seconds per instance for all CWRU load conditions, and graph construction consistently takes less than 0.14 seconds. Although the SU dataset exhibits higher feature extraction time due to its greater number of channels and more complex signals (e.g., 23.05 seconds for the Bearing set), the inference stage remains in the sub-millisecond range (below 0.001 s), confirming the model’s scalability and real-world viability.
Collectively, these results position the proposed model not only as a state-of-the-art solution in diagnostic accuracy but also as a highly efficient and deployable system for industrial prognostics and health management.
4 Discussion
In this framework, we evaluate three classical classifiers, logistic regression (LR), random forests (RF), and support vector machines (SVM), on features extracted from the benchmark datasets. Based on comparative performance, random forests emerged as the most effective classifier, offering superior accuracy and robustness under various conditions.
Based on the evaluation results, the random forest classifier was selected owing to its consistently superior performance. The proposed graph-based fault detection framework demonstrates robust effectiveness across both the CWRU and SU datasets, effectively handling a wide range of operating conditions and fault types. On the CWRU dataset, the model achieved high accuracy and F1-scores under various machine loads, with the best performance observed at lower loads (0 and 1 HP), where amplified fault signatures enhanced both the classification accuracy and graph structure clarity. In contrast, the performance slightly declined at higher loads (2 and 3 HP), likely because of less distinctive fault signals. In the SU dataset, which features controlled and balanced conditions, the framework achieved near-perfect classification across all fault types, highlighting its sensitivity to subtle variations in vibration patterns and its adaptability across datasets with differing characteristics
However, the CWRU dataset poses limitations in terms of generalizability. Its idealized laboratory conditions, artificially induced faults via controlled EDM cuts, minimal ambient noise, and fixed load variations fail to replicate the natural degradation, signal interference, or dynamic loads typical of industrial environments. To address this, we incorporated the SU gearbox dataset, which includes more realistic noise, load conditions, and fault combinations, providing a stronger foundation for assessing real-world applicability.
Transfer learning experiments confirmed near-perfect classification under same-load conditions; however, cross-load performance slightly declined with increasing load disparities, as validated by statistical analysis. In complex scenarios, including cross-fault, cross-component, and combined load-and-fault conditions, Random Forest classifiers outperformed logistic regression, particularly in noisy or imbalanced settings, owing to their superior handling of variability. Graph construction parameters significantly influenced the performance, such as $k_{\text{max}}$ and subgraph size. Mid-sized subgraphs exhibiting small-world properties and higher modularity optimize fault detection accuracy while maintaining computational efficiency. Noise analysis further underscored the robustness of the model, with stable accuracy in the SU dataset under high noise levels.
SHAP values quantify feature contributions to random forest decisions. Local time-frequency features (e.g., RMS, skewness, kurtosis) dominated predictive power, while graph-theoretic metrics such as modularity (Q), spectral gap, and average path lengthoffered meaningful, albeit smaller, contributions. These metrics capture complementary signal aspects: modularity reflects fault-specific community cohesion, the spectral gap indicates network robustness across fault severities, and the average path length measures information propagation delays from fault patterns. Together, they improve the separation of subtle or overlapping fault classes beyond local statistics alone.
Adaptive segmentation optimizes the window and step sizes by maximizing a composite entropy score (encompassing amplitude, envelope, spectral, and envelope-spectrum entropies). High-entropy windows align with informative transient fault events, whereas low-entropy windows often represent noise or steady states. This entropy maximization suppresses noisy segments and prioritizes fault-relevant bursts, thereby improving the segment-level feature quality. Under synthetic Gaussian noise ($\sigma=0.5$), adaptive segmentation sustained ¿90% accuracy on the CWRU dataset under 0 and 1 HP loads (¿85% otherwise) and ¿95% on the SU dataset, illustrating its adaptability to signal dynamics and robust noise tolerance for enhanced graph representations.
Our framework offers notable advantages over deep learning methods such as Graph Convolutional Networks (GCNs), spatial-temporal graphs, and graph autoencoders. These methods often require extensive computational resources, lack interpretability, and are prone to dataset biases, hindering practical deployment. Our approach, which leverages entropy-based adaptive segmentation, explicit time-frequency features, and interpretable graph metrics (e.g., modularity, spectral gap, and average path length), ensures robust generalization, computational efficiency, and transparency. Integrating feature-rich graph representations with simple classifiers, such as random forests, makes the model accessible and well-suited for real-time or embedded industrial systems, where explainability and reliability are paramount. Interpretable graph metrics enable engineers to trace vibration anomalies to specific fault types, thereby enhancing diagnostic precision and practical utility.
However, challenges persist for real-world deployment. The model’s dependence on finely tuned segmentation and graph construction parameters may require adjustments for machines with different vibration profiles. Although robust across the tested conditions (load, fault type, and noise), adapting to new sensor configurations, mixed-signal sources, or dynamic system behaviors may require further refinements. Preprocessing steps, such as time-frequency feature extraction and graph creation, can introduce delays in time-sensitive industrial settings, necessitating hardware acceleration or parallel processing. Additionally, the model’s performance relies on adequate data length and quality; short signals, missing data, or irregular sampling rates may compromise its effectiveness.
Future work should focus on online or continual learning to enable real-time adaptation with minimal manual intervention and ensure compatibility with dynamic industrial environments. Advanced feature engineering can enhance performance while preserving interpretability and efficiency. Exploring domain generalization strategies, such as adversarial training or self-supervised learning, could improve robustness across diverse machine types and operating conditions without retraining. These advancements bolster the framework’s scalability and autonomy, making it highly suitable for complex industrial applications.